OpenStudy (anonymous):

A number, twice that number, and one third of that number are added. The result is 20. What is the original number?

OpenStudy (anonymous):

Let the number be "x". EQUATION: x + 2x + (1/3)x = 20 (10/3)x = 20 (1/3)x = 2 x = 6-------- answer

OpenStudy (anonymous):

ur welcome.